OpenStudy (anonymous):

Find the equation of the line that passes through the points (2, 4) and (4, 2).

OpenStudy (texaschic101):

first we need to find the slope by using the slope formula. slope(m) = (y2 - y1) / (x2 - x1) (2,4) x1 = 2 and y1 = 4 (4,2) x2 = 4 and y2 = 2 now sub in our info... m = (2 - 4) / (4 - 2) m = -2/2 m = -1 <-- your slope you with me so far ?

OpenStudy (texaschic101):

Now that we know the slope, and we have points, we can use the point slope formula. y - y1 = m(x - x1) m (slope) = -1 and we can use either of the points, I will use (2,4) (2,4) x1 = 2 and y1 = 4 now we will sub y - 4 = -1(x - 2) y - 4 = -x + 2 y = -x + 2 + 4 y = -x + 6 <-- your answer any questions ?